# Imagine to Hear: Auditory Knowledge Generation can be an Effective Assistant for Language Models

### v1

@@ Abstract
Language models pretrained on text-only corpora often struggle with tasks that require auditory commonsense knowledge. Previous work addresses this problem by augmenting the language model to retrieve knowledge from external audio databases. This approach has several limitations, such as the potential lack of relevant audio in databases and the high costs associated with constructing and querying the databases. To address these issues, we propose Imagine to Hear, a novel approach that dynamically generates auditory knowledge using generative models. Our framework detects multiple audio-related textual spans from the given prompt and generates corresponding auditory knowledge. We develop several mechanisms to efficiently process multiple auditory knowledge, including a CLAP-based rejection sampler and a language-audio fusion module. Our experiments show that our method achieves state-of-the-art performance on AuditoryBench without relying on external databases, highlighting the effectiveness of our generation-based approach.
@@ 1 Introduction
While such limitations can be partially addressed by augmenting language models with auditory representations retrieved from external audio database (Ok et al., 2025), this approach suffers from two key limitations: (1) The database may not contain any audio that is directly relevant to the given query. (2) Constructing and querying such databases requires a large computational cost.
@@ 3 Experiments
We conducted our experiments on the AuditoryBench with further experiments. Details provided in Appendix C.
@@
To evaluate the effectiveness of our approach, we compare it with recent studies Ok et al. (2025) on AuditoryBench. As shown in Table 1, most language models perform poorly due to the absence of auditory knowledge. While AudioBERT performs well by leveraging external databases, it struggles with inter-domain, as reflected in the wiki test results. In contrast, our method leverages imagination to dynamically generate relevant auditory representations, establishing a new state-of-the-art AuditoryBench. This result highlights the enhanced generalization ability of our approach.
@@
We first examine the effect of removing the rejection system. As shown in Figure 2 (a), the rejection threshold of 0 is optimal for animal sound recognition. In contrast, for sound pitch comparison, the optimal threshold is 0.6, as illustrated in Figure 2 (b). For sound pitch comparison, to further analyze the impact of the iteration count $n$, we conducted an additional ablation study by fixing the rejection threshold at 0.6 (see Figure 2 (c)). The results indicate that accuracy improves as $n$ increases, reaching its peak at $n=2$. Detailed performance improvements are reported in Table 2.
@@
Role of fusion gate
@@
Our analysis shows that the fusion gate effectively prioritizes auditory-relevant tokens while minimizing the influence of function words. In phrases like "ambient noise of a living room" and "sound of a household draining system," key auditory descriptors such as "ambient," "living," "room," "household," "draining," and "system" receive the highest weights. This indicates that the model assigns more importance to tokens that directly describe sound characteristics while suppressing function words like "of" and "a", which contribute less to auditory integration.
@@
We also illustrate more ablation studies and case studies for each method to demonstrate their contributions in Appendix B.
@@ 4 Conclusion
One limitation is that the effectiveness of our method is influenced by the quality of input prompts. If the input lacks sufficient auditory cues, the generated audio may not always align perfectly with the intended knowledge (Appendix D). Additionally, while our rejection-based generation ensures high-quality auditory integration, it introduces some computational overhead.
@@ Appendix A Related work
Appendix B More ablation study and case studies
In this section, we describe more ablation studies and case studies of our methods.
Table 3: diffusion ablation
@@ Appendix B More ablation study and case studies
B.1 Various generation models.
@@
B.2 Case Study
To further illustrate the effectiveness of our approach, we visualize more case studies that our method efficiently works.
Dynamic knowledge injection
To further investigate the role of dynamic knowledge injection (DKI), we analyze cases where injecting auditory knowledge at the span level improves model predictions, as illustrated in Figure 5.
In the first example, the sentence compares the sound of a guitar with that of a sopranissimo saxophone. When generating audio representations separately for each instrument, their mel-spectrograms exhibit distinct frequency characteristics, with the saxophone producing significantly higher frequencies than the guitar. However, when generating a single audio representation for the entire sentence, this distinction is lost, leading to an incorrect prediction of "higher." In contrast, our method, by dynamically injecting span-specific auditory knowledge, correctly predicts "lower," maintaining the frequency contrast observed in the spectrograms.
In the second case, the sentence contrasts the sound of an alto trombone with that of a sopranino trombone. Similarly, generating audio separately for each instrument produces clear frequency differences, with the sopranino trombone having higher frequencies. However, sentence-level generation fails to capture this difference, leading to an incorrect prediction of "higher." By selectively injecting auditory spans, our model correctly predicts "lower," aligning with the frequency distribution observed in the mel-spectrogram analysis.
Table 4: Statistics for Auditorybench.
@@ B.2 Case Study
Appendix C Experiment details
C.1 Training Datasets
@@
The dataset includes 6,212 sentences for Animal Sound Recognition and 15,502 sentences for Sound Pitch Comparison, with varying lengths and complexities. The average number of words per sentence is 9.21 for Animal Sound Recognition and 16.83 for Sound Pitch Comparison. More statics info is in Table 4.
@@
Appendix D Experiments on GLUE
@@
To detect auditory spans within GLUE, we train an audio span detector based on a BERT-base model. The detector is trained for 5 epochs with a batch size of 16, a learning rate of $1\times 10^{-5}$, and the AdamW optimizer Loshchilov. and Hutter (2019). Since GLUE lacks predefined auditory span labels, we generate labels using Qwen2-72B-Instruct-AWQ Yang et al. (2024) that can be inferred on a single A100 GPU. The model independently generates labels for auditory spans using a predefined prompt, which is detailed in Appendix F.
Appendix E Detailed dataset information of GLUE
@@
Appendix F Prompts for span generation

### v2

@@ Abstract
Language models pretrained on text-only corpora often struggle with tasks that require auditory commonsense knowledge. Previous work addresses this problem by augmenting the language model to retrieve knowledge from external audio databases. This approach has several limitations, such as the potential lack of relevant audio in databases and the high costs associated with constructing the databases. To address these issues, we propose Imagine to Hear, a novel approach that dynamically generates auditory knowledge using generative models. Our framework detects multiple audio-related textual spans from the given prompt and generates corresponding auditory knowledge. We develop several mechanisms to efficiently process multiple auditory knowledge, including a CLAP-based rejection sampler and a language-audio fusion module. Our experiments show that our method achieves state-of-the-art performance on AuditoryBench without relying on external databases, highlighting the effectiveness of our generation-based approach.
@@ 1 Introduction
While such limitations can be partially addressed by augmenting language models with auditory representations retrieved from external audio database (Ok et al., 2025), this approach suffers from two key limitations: (1) The database may not contain any audio that is directly relevant to the given query. (2) Constructing such databases requires a large computational cost.
@@ 3 Experiments
We conducted our experiments on the AuditoryBench with further experiments. More details are provided in Appendix D.
@@
Task Dev. Test Wiki
Animal sound 92.47 $\pm$ 0.77 92.66 $\pm$ 0.35 100.00 $\pm$ 0.00
Sound pitch 94.50 $\pm$ 0.32 94.75 $\pm$ 0.29 90.24 $\pm$ 3.00
Table 2: F1 scores of span detection.
@@
To evaluate the effectiveness of our approach, we compare it with recent studies Ok et al. (2025) on AuditoryBench. As shown in Table 1, most language models perform poorly due to the absence of auditory knowledge. While AudioBERT performs well by leveraging external databases, it struggles with inter-domain, as reflected in the wiki test results. In contrast, our method leverages imagination to dynamically generate relevant auditory representations, establishing a new state-of-the-art AuditoryBench. This result highlights the enhanced generalization ability of our approach. We additionally report F1 scores of our span detector, in Table 2. We observe that even a simple span detector can achieve high accuracy across all subsets.
@@
We first examine the effect of removing the rejection system. As shown in Figure 2 (a), the rejection threshold of 0 is optimal for animal sound recognition. In contrast, for sound pitch comparison, the optimal threshold is 0.6, as illustrated in Figure 2 (b). For sound pitch comparison, to further analyze the impact of the iteration count $n$, we conducted an additional ablation study by fixing the rejection threshold at 0.6 (see Figure 2 (c)). The results indicate that accuracy improves as $n$ increases, reaching its peak at $n=2$. Detailed performance gains are reported in Table 3.
@@
Role of fusion gate.
@@
Our analysis shows that the fusion gate effectively prioritizes auditory-relevant tokens while minimizing the influence of function words. In phrases like "ambient noise of a living room" and "sound of a household draining system," key auditory descriptors such as "ambient," "living," "room," "household," "draining," and "system" receive the highest weights. This indicates that the model assigns more importance to tokens that directly describe sound characteristics while suppressing function words like "of" and "a," which contribute less to auditory integration.
Dynamic knowledge injection.
To deepen our understanding on the role of dynamic knowledge injection (DKI), we analyze cases where injecting auditory knowledge at the span level improves model predictions, as illustrated in Figure 5.
In the first example, the sentence compares the sound of a guitar with that of a sopranissimo saxophone. When generating audio representations separately for each instrument, their mel-spectrograms exhibit distinct frequency characteristics, with the saxophone producing significantly higher frequencies than the guitar. However, when generating a single audio representation for the entire sentence, this distinction is lost, leading to an incorrect prediction of “higher.” In contrast, our method, by dynamically injecting span-specific auditory knowledge, correctly predicts “lower,” maintaining the frequency contrast observed in the spectrograms.
In the second case, the sentence contrasts the sound of an alto trombone with that of a sopranino trombone. Similarly, generating audio separately for each instrument produces clear frequency differences, with the sopranino trombone having higher frequencies. However, sentence-level generation fails to capture this difference, leading to an incorrect prediction of “higher.” By selectively injecting auditory spans, our model correctly predicts “lower,” aligning with the frequency distribution observed in the mel-spectrogram analysis.
@@
We also illustrate more ablation studies for each method to demonstrate their contributions in Appendix B.
@@ 4 Conclusion
ITH extends beyond its established advantages over retrieval methods in the multimodal era. Its utility is notable in resource-constrained settings reliant on unimodal models where multimodal inputs are infeasible. Furthermore, even for advanced multimodal systems, ITH can provide a practical pathway when audio is absent, ambiguous, or subtle—parallel to visual imagination modules that enhance reasoning without direct visual input (Lu et al., 2022; Zhu et al., 2023; Liu et al., 2024).We expect that our approach will remain effective and complementary in the era of multimodal language models.
@@
One limitation is that the effectiveness of our method is influenced by the quality of input prompts. If the input lacks sufficient auditory cues, the generated audio may not always align perfectly with the intended knowledge (Appendix E). Additionally, while our rejection-based generation ensures high-quality auditory integration, it introduces some computational overhead.
Acknowledgments
This work was supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Ministry of Science and ICT (MSIT)(No.RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH)), and in part by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT)(No. RS-2023-00213710, neural network optimization with minimal optimization costs). We also thank our colleagues at EffL Lab and others for their help.
@@ Appendix A Related work
Appendix B More ablation study
@@
Table 4: diffusion ablation
Various generation models.
@@
Appendix C Human Evaluation
Generated Audio.
To assess perceptual quality of the generated audio samples, we have conducted human evaluations comparing our generated audio with real audio retrieved from LAION-Audio-630K. Sixteen participants have evaluated the samples under both in pairwise and 3-way identification settings, with the total of 40 questions. The results, shown in Table 5 and Table 6, indicate that the generated audio is often indistinguishable from the real recordings. Specifically, Table 5 shows that, when comparing retrieval-based audio (in AudioBERT) and generated audio (in ours) conditioned on the same input text from AuditoryBench, human evaluators exhibit a clear preference for the generated audio. Table 6 presents a 3-way identification task where participants are asked to identify the real sample among two generated and one real audio; the results suggest that participants struggle to distinguish real from generated audio reliably.
Preference Ratio (%)
Real Audio 41.16
Generated Audio 50.30
Equal 8.54
Table 5: Pairwise comparison between real and generated audio.
Setup Chose Real (%) Chose Generated (%)
2 Reals, 1 Gen 65.00 35.00
2 Gens, 1 Real 44.03 55.97
Table 6: We design a 3-way identification task to evaluate the realism of generated audio. In the first setup, participants are presented with two real audio samples and one generated sample and are asked to identify which one is the generated audio. In the second setup, the inverse task is given: among the three audio samples (two generated and one real), participants are asked to identify the real audio.
We have further evaluated on the wiki set, which includes real audio aligned with the prompts. As shown in Table 7, ten out of sixteen annotators have judged the generated audio to be of sufficiently high quality and fidelity to serve as plausible substitutes; we have tested with 20 questions sampled from the wiki set. These findings confirm that our method produces realistic and contextually appropriate auditory knowledge.
Evaluation Criterion Result
Acceptable as substitute for real audio 77.42%
Quality and semantic fidelity (1–5 scale) 3.78
Table 7: Human evaluation on the Wikipedia subset of AuditoryBench.
@@ Appendix C Human Evaluation
Table 11: Statistics for GLUE benchmark datasets.
Appendix D Experiment details
D.1 Training Datasets
@@
The dataset includes 6,212 sentences for Animal Sound Recognition and 15,502 sentences for Sound Pitch Comparison, with varying lengths and complexities. The average number of words per sentence is 9.21 for Animal Sound Recognition and 16.83 for Sound Pitch Comparison. More statistics info is in Table 8.
@@
Appendix E Experiments on GLUE
@@
To detect auditory spans within GLUE, we train an audio span detector based on a BERT-base model. The detector is trained for 5 epochs with a batch size of 16, a learning rate of $1\times 10^{-5}$, and the AdamW optimizer Loshchilov. and Hutter (2019). Since GLUE lacks predefined auditory span labels, we generate labels using Qwen2-72B-Instruct-AWQ Yang et al. (2024) that can be inferred on a single A100 GPU. The model independently generates labels for auditory spans using a predefined prompt, which is detailed in Appendix G.
Appendix F Detailed dataset information of GLUE
@@
Appendix G Prompts for span generation